Suppose all the electrons of 100 g water are lumped together to form a negatively charged particle and all the nuclei are lumped together to form a positively charged particle. If these two particle are placed 10.0 cm away from each other, find the force of attraction between them. Compare it with your weight.

Grade:upto college level

2 Answers

Deepak Patra
9 years ago
Sol. Molecular weight of H base 2O = 2 × 1 × 16 = 16 No. of electrons present in one molecule of H base 2O = 10 18 gm of H base 2O has 6.023 × 10^23 molecule 18 gm of H base 2O has 6.023 × 10^23 × 10 electrons 100 gm of H base 2O has 6.023 * 10^24/18 * 100 elecrons So number of protons = 6.023 * 10^26/18 protons (since atom is electrically neutral) Charge of protons = 1.6 * 10^-19 * 6.023 * 10^26/18 coulomb = 1.6 * 6.023 * 10^7/18 coulomb Charge of electrons = = 1.6 * 6.023 * 10^7/18 coulomb Hence Electrical force = 9 * 10^9(1.6 * 6.023 * 10^7/18)*( 1.6 * 6.023 * 10^7/18)/(10 * 10^-2)^2 = 8 * 6.023/18 * 1.6 * 6.023 * 10^25 = 2.56 * 10^25 Newton
maha
7 years ago
Suppose all the electrons of 100 g water are lumped together to form a negatively charged particle and all the nuclei are lumped together to form a positively charged particle. If these two particle are placed 10.0 cm away from each other, find the force of attraction between them. Compare it with your weight.